### v5

# The formation of Schrödinger cat-like states in the process of spontaneous parametric down-conversion

Abstract
We investigate the formation of entangled Schrödinger cat-like states (SCLSs) in both the fundamental and second-harmonic modes during spontaneous parametric down-conversion (SPDC), considering quantum depletion of the pump. Unlike conventional semiclassical treatments that assume an undepleted classical pump, we fully quantize both modes and explore their dynamics under non-dissipative and dissipative regimes. Using the Wigner function, we visualize the non-Gaussian character and interference patterns indicative of these entangled states. Our analysis shows significant squeezing (up to $3.76\,\mathrm{dB}$ without dissipation and $3.65\,\mathrm{dB}$ with dissipation) and super-Poissonian statistics (Fano factor $\gg 1$), both persisting under moderate losses—features that extend beyond those of conventional optical Schrödinger cat states, which lack such robustness. In contrast to semiclassical models yielding statistical mixtures, our approach captures energy exchange between modes, phase-dependent evolution, interference, and interaction-length sensitivity. These features are critical for continuous-variable quantum technologies. The generated entangled SCLSs are highly relevant for applications in quantum sensing and quantum information processing. Our findings pave the way for realizing strongly non-Gaussian entangled states using $\chi^{(2)}$ media at moderate pump intensities, offering advantages over $\chi^{(3)}$-based schemes.
preprint: APS/123-QED
I Introduction
We investigate the formation of entangled Schrödinger cat-like states (SCLSs) in the fundamental and second-harmonic modes of spontaneous parametric down-conversion (SPDC) under a fully quantized pump model. Unlike conventional optical Schrödinger cat states (even/odd coherent states [Dodonov1974]) that exhibit only non-classical interference, our SCLSs demonstrate additional quantum features including significant squeezing (up to $3.76\,\mathrm{dB}$) and super-Poissonian statistics (Fano factor $\gg 1$), both robust under moderate losses. These properties, combined with their inherent entanglement between modes, distinguish them fundamentally from single-mode cat states. Going beyond undepleted-pump approximations, we analyze their dynamics in non-dissipative and dissipative regimes. Wigner function visualizations reveal the non-Gaussian character and interference patterns of these states, confirming their SCLS nature. Unlike semiclassical models that produce statistical mixtures, our approach captures energy exchange, phase-dependent evolution, and interference, highlighting critical properties for continuous-variable quantum technologies. These entangled SCLSs are promising for quantum sensing and information processing, offering a path to strong non-Gaussianity in $\chi^{(2)}$ media at moderate pump intensities.
While the standard treatment of SPDC approximates the Hamiltonian as quadratic by treating the pump classically, our work employs the full cubic Hamiltonian $\hat{H}_{\mathrm{int}}\propto\hat{a}_{1}^{\dagger 2}\hat{a}_{2}+\mathrm{h.c.}$ with a quantized pump. The key novelty lies in exploring the high-depletion operating condition—where specific initial states lead to approximately 64% pump conversion—whose quantum features had remained largely uncharacterized due to the dominance of the undepleted-pump approximation. This operating condition could be realized with modern highly-efficient nonlinear crystals (e.g., periodically poled structures [see Belinsky2018, and references therein]) and using contemporary numerical tools like QuTiP [johansson2012qutip, JOHANSSON20131234].
Schrödinger cat-like states, characterized by negative values of the Wigner function, play an important role in, e.g., observing phase displacements due to their phase sensitivity [Shukla2023, Toscano2006, Salykina2023, Singh2024]. Schrödinger cat states can be used to encode cat qubits [Cochrane1999] and building Ising machines [Yamamoto2020].
In the field of quantum nonlinear optics, the generation of non-Gaussian states requires at least the cubic form of the interaction Hamiltonian [Braunstein2005]. The formation of non-Gaussian states [Sychev2017] (e.g., Schrödinger cat-like states) has been considered in [Nikitin1991, Singh2021] during quantum state evolution in second harmonic generation with $\chi^{(2)}$ media, with [Singh2021] demonstrating interference patterns using Wigner functions.
The classical and non-depleted or semi-classical method approximates the interaction Hamiltonian for the parametric down-conversion (PDC) process, which is based on $\chi^{(2)}$ (and is cubic), by a quadratic one (i.e., pump mode is considered classic and undepleted). Such an approximation neglects the non-Gaussianity features present in the initial cubic form of the interaction Hamiltonian. Another approximation method, e.g., the expansion of unitary operators [Agarwal1974, Beskrovnyi1996, Belinsky2018], can be applied to consider a full treatment of quantum effects present in all the modes of the PDC process, but starts to deviate at later stages [Belinsky2018]. However, the diagonalization method [Walls1970, Walls1972, Gantsog1991, Tanas1992, Belinsky2018] can be used to study evolution at longer interaction lengths without losing the quantum effects present in the interacting modes.
Previous studies have implemented a full quantum mechanical treatment of parametric down-conversion (PDC) with a depleted pump mode [Walls1970, Walls1972, Gantsog1991, Tanas1991, Tanas1992, Drobny1992], investigating statistical properties including mean photon numbers, quadrature fluctuations in the fundamental mode, and phase properties of both modes [Gantsog1993].
Conversely, semiclassical treatments [w1988, Walls2008] assume a classical, undepleted pump, leading to a Fokker-Planck equation whose steady-state solution yields a statistical mixture of two coherent states. This mixture inherently lacks crucial quantum features: it exhibits no interference pattern, cannot form genuine even or odd cat states, omits energy exchange between modes, shows no dependence on interaction length, and is confined to the dissipative regime.
In contrast, our work demonstrates the formation of SCLSs in both modes across both non-dissipative and dissipative regimes. Our results demonstrate the preservation of the quantum interference structure and provide clear evidence of energy exchange between the fundamental and second-harmonic modes (see Fig. 1 and Fig. 2).
In this paper, we have studied the formation of SCLSs for two cases (non-dissipative and dissipative) when both modes (fundamental and second harmonic) are considered as quantum and depleted during the effective realization of the SPDC process based on $\chi^{(2)}$ [Walls2008, Agarwal2013]. QuTiP [johansson2012qutip, JOHANSSON20131234] is used to numerically solve the Lindbladian superoperator responsible for the SPDC process. The formation of SCLSs is studied and illustrated qualitatively using the Wigner function. Quantitative analysis of the SCLSs is performed by calculating squeezing levels (variances of quadrature components and Fano factor), values of photon number distributions and the entanglement between the two modes. The SCLSs are observed in the fundamental and second harmonic modes for non-dissipative and dissipative cases when the fundamental mode is in the vacuum state and the second harmonic mode is in the coherent state.
II SPDC under pump depletion and losses
Let three stationary degenerate monochromatic optical plane wave modes $\hat{a}_{s},\hat{a}_{i},\hat{a}_{p}$ of frequencies $\omega_{s}$, $\omega_{i}$, $\omega_{p}$ propagate collinearly in a nonlinear optical crystal with non-zero susceptibility $\chi^{(2)}$. Subscripts $i,s,p$ belong to the idler, signal, and pump modes. In case of degenerate frequencies: $\omega_{s}=\omega_{i}=\omega$, $\omega_{p}=2\omega$, $\hat{a}_{i}=\hat{a}_{s}=\hat{a}_{1},\hat{a}_{p}=\hat{a}_{2}$. For effective realization of the SPDC process, it is assumed that all three modes can be phase-matched [Dmitriev1999] or quasi-phase-matched [Belinsky2018]. The interaction Hamiltonian of the SPDC process $(2\omega=\omega+\omega)$ [Walls2008, Agarwal2013] is given by
$\hat{H}_{int}=\hbar g(\hat{a}^{2}_{1}\hat{a}_{2}^{\dagger}+h.c.),$ (1)
where $\hbar$ is the Planck constant. For simplicity $\hbar=1$. $g$ is the nonlinear coupling constant of the interacting modes.
The Lindblad master equation for the density matrix $\hat{\rho}$ describing an open quantum system is used to study the quantum dynamics of the interaction Hamiltonian (1 and can be written as [Agarwal2013, Walls2008, BP2002, Strinati2024]
[…]
where $\tau=gt$ is the normalized interaction length. $\hat{C}_{j}=\sqrt{\gamma_{j}/g}\hat{a}_{i}$ are the Lindblad operators describing the dissipative part of the dynamics, where $\gamma_{j}\geq 0$ are the cavity damping rates of the modes and $g>0$.
Equation (2) is numerically solved by using QuTiP for two cases (non-dissipative $\gamma_{j}=0$ and dissipative $\gamma_{j}=0.15$) for the initial state density matrix $\hat{\rho}(0)=|\psi_{0}\rangle\langle\psi_{0}|$. At the input of the nonlinear crystal, the fundamental mode is in the vacuum state $|0\rangle_{1}$, the second harmonic one is in the coherent state $|\alpha_{20}\rangle_{2}=e^{-|\alpha_{20}|^{2}/2}\sum_{n_{2}=0}^{\infty}\frac{\alpha_{20}^{n_{2}}}{\sqrt{n_{2}!}}|n_{2}\rangle$, having mean number of photons $|\alpha_{20}|^{2}=20$ and phase $\varphi_{20}=\pi/2$, i.e., $|\psi_{0}\rangle=|0\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$.
III Formation of cat-like states in both modes
One can study and analyze the quantum statistical properties of the modes by using the Wigner quasiprobability distribution [Walls2008, Agarwal2013]. Phase space portraits of the Wigner function help to visualize delicate patterns, such as (a) interference (wave nature) present in the superposition of macroscopically distinct states, e.g., SCLSs, (b) non-Gaussianity of the state, i.e., the negative value of the Wigner function. The Wigner functions of the modes $\hat{a}_{1}$ and $\hat{a}_{2}$ are calculated using
$W_{j}(\alpha_{j},\tau)=\frac{1}{\pi}\int_{-\infty}^{\infty}\langle\Re{(\alpha_{j})}-y|\hat{\rho}_{j}(\tau)|\Re{(\alpha_{j})}+y\rangle e^{2i\Im{(\alpha_{j})}y}dy,$ (3)
where $j=1,2$, $\hat{\rho}_{1}(\tau)=\operatorname{Tr}_{2}[\hat{\rho}(\tau)]$, $\hat{\rho}_{2}(\tau)=\operatorname{Tr}_{1}[\hat{\rho}(\tau)]$ are reduced density matrices of states of the modes $\hat{a}_{1}$ and $\hat{a}_{2}$.
Figure 1: Wigner functions $W_{1}(\alpha_{1},\tau)$ and $W_{2}(\alpha_{2},\tau)$ during SPDC at $\tau=0.38$. Top row (a,b): Non-dissipative ($\gamma_{1,2}=0$), $N_{1}\approx 25.2$, $N_{2}\approx 7.36$. Bottom row (c,d): Dissipative ($\gamma_{1,2}=0.15$), $N_{1}\approx 23.88$, $N_{2}\approx 6.95$. Initial: $\hat{a}_{1}$ vacuum, $\hat{a}_{2}$ coherent state ($|\alpha_{20}|^{2}=20$, $\varphi_{20}=\pi/2$).
In Fig. 1 plots of (3) are shown for two cases (non-dissipative and dissipative) of the formation of SCLSs in the fundamental and second harmonic modes.
IV Application-relevant statistical properties
Using the Wigner function, we calculated the contribution of all statistical moments present in the SCLSs and qualitatively visualized portraits of the Wigner functions in phase space (see Fig. 1). In order to quantitatively estimate quantum statistical properties such as squeezing level (the variances of quadrature components, the Fano factor) photon number distribution calculations and the entanglement measure are performed.
IV.1 Mean number extrema as indicators of SCLSs formation. Squeezing in both modes
The mean number of photons in the both modes $\hat{a}_{j}$ is calculated using
$N_{j}(\tau)=\operatorname{Tr}[\hat{a}^{\dagger}_{j}\hat{a}_{j}\hat{\rho}_{j}(\tau)].$ (4)
The variances of the quadrature components are calculated using
$\Delta^{2}x_{j}=\operatorname{Tr}[\hat{x}_{j}^{2}\hat{\rho}_{j}(\tau)]-\operatorname{Tr}[\hat{x}_{j}\hat{\rho}_{j}(\tau)]^{2},$ (5)
$\Delta^{2}p_{j}=\operatorname{Tr}[\hat{p}_{j}^{2}\hat{\rho}_{j}(\tau)]-\operatorname{Tr}[\hat{p}_{j}\hat{\rho}_{j}(\tau)]^{2},$ (6)
where $\hat{x}_{j}=2^{-(1/2)}(\hat{a}_{j}+\hat{a}^{\dagger}_{j})$ and $\hat{p}_{j}=-i2^{-(1/2)}(\hat{a}_{j}-\hat{a}^{\dagger}_{j})$ are the quadrature components of the modes $\hat{a}_{j}$. Fig. 2 shows the evolution of both (i) the mean photon numbers (Eq. 4) and (ii) the quadrature component variances (Eqs. 5 and 6) for the non-dissipative case. SCLSs form at the first extremal values of the photon numbers: the maximum for mode $\hat{a}_{1}$ and the minimum for mode $\hat{a}_{2}$, both occurring at $\tau=0.38$. Notably, the quadrature variances $\Delta^{2}p_{1}(0.38)=0.23$ (mode $\hat{a}_{1}$) and $\Delta^{2}x_{2}(0.38)=0.21$ (mode $\hat{a}_{2}$) demonstrate squeezing, corresponding to 3.37-3.37 dB and 3.76-3.76 dB respectively.
Figure 2: Evolution of (a) $N_{j}$ and (b) quadrature variances SPDC process. $N_{1}$ ($\hat{a}_{1}$, solid black, -) peaks while $N_{2}$ ($\hat{a}_{2}$, dashed black, --) minimizes at $\tau=0.38$. Variances: $\Delta^{2}x_{1}$ (blue $\cdot{\cdot}{\cdot}$), $\Delta^{2}p_{1}$ (blue $\cdot$-), $\Delta^{2}x_{2}$ (red $\cdot{\cdot}{\cdot}$), $\Delta^{2}p_{2}$ (red $\cdot$-). $\bigstar$ marks extrema.
Figure 3: Photon number distributions $P_{j}(n_{j})$ at $\tau=0.38$: (a,b) non-dissipative ($\gamma_{1,2}=0$) and (c,d) dissipative ($\gamma_{1,2}=0.15$) regimes for modes $\hat{a}_{1}$ (a,c) and $\hat{a}_{2}$ (b,d).
IV.2 Dissipation-induced parity mixing in SCLSs
It helps to identify the type of pattern an SCLS belongs to, e.g. even, odd, mixture of even and odd coherent states. The photon number distribution is calculated using
$P_{j}(n_{j})=\operatorname{Tr}\left[\hat{\rho}_{j}{|n_{j}\rangle}{\langle n_{j}|}\right].$ (7)
Fig. 3 shows pattern types of photon number distributions (7) of SCLSs $\hat{\rho}_{j}$. In case of non-dissipative regime, photon number distributions of SCLSs $\hat{\rho}_{j}$ for modes $\hat{a}_{1}$ and $\hat{a}_{2}$ are associated with even and mixture of even and odd coherent states. In the case of the dissipative regime, Figs. 3(c)-3(d) show that the SCLSs $\hat{\rho}_{j}(\tau)$ are associated with the photon number distribution of the mixture of even and odd coherent states. The dissipative regime introduces an odd number of photons into the SCLSs $\hat{\rho}_{j}(\tau)$ (see Figs. 3(c)-3(d)).
IV.3 Robustness of super-Poissonian statistics to dissipation
The value of the Fano factor can identify the type (sub-Poissonian, Poissonian, super-Poissonian) of the distributions of the studied SCLS $\hat{\rho}_{j}$. Thus, the Fano factor [Vogel] is calculated using
$FF_{j}(\tau)=\frac{\operatorname{Tr}[(\hat{a}^{\dagger}_{j}\hat{a}_{j})^{2}\hat{\rho}_{j}(\tau)]-N_{j}^{2}(\tau)}{N_{j}(\tau)}.$ (8)
The Fano factor $FF_{j}$ characterizes photon statistics: $FF_{j}<1$ indicates sub-Poissonian, $FF_{j}=1$ Poissonian, and $FF_{j}>1$ super-Poissonian behavior. At $\tau=0.38$, the calculated values from Eq. (8) are: $FF_{1}=8.54$ (non-diss.), $8.41$ (diss.); $FF_{2}=6.67$ (non-diss.), $6.48$ (diss.)—all indicating super-Poissonian statistics.
IV.4 Entanglement of SCLSs
To quantify the entanglement between the two modes $(\hat{a}_{1},\hat{a}_{2})$ in the SCLSs, we compute the Schmidt number [Law2004] defined as:
$K(\tau)=\frac{1}{\operatorname{Tr}[(\hat{\rho}_{1}(\tau))^{2}]}=\frac{1}{\operatorname{Tr}[(\hat{\rho}_{2}(\tau))^{2}]},$ (9)
where $\hat{\rho}_{1}$ and $\hat{\rho}_{2}$ are the reduced density matrices of modes $\hat{a}_{1}$ and $\hat{a}_{2}$, respectively. For pure bipartite states, the parameter $K$ serves as an entanglement measure, with $K>1$ indicating non-separable (entangled) states. At $\tau=0.38$, the value calculated from Eq. (9) is $K=1.93$ for the non-dissipative case, confirming the presence of entanglement.
V Summary and conclusions
We demonstrate the formation of entangled Schrödinger cat-like states (SCLSs) in both fundamental and second-harmonic modes during spontaneous parametric down-conversion (SPDC) with quantum-depleted pumps, under both non-dissipative ($\gamma_{1,2}=0$) and dissipative ($\gamma_{1,2}=0.15$) regimes. At $\tau=0.38$, Wigner functions reveal clear SCLS formation with mode $\hat{a}_{1}$ exhibiting $N_{1}\approx 25.2$ and mode $\hat{a}_{2}$ showing $N_{2}\approx 7.3$ in the non-dissipative case, while maintaining non-Gaussianity (Wigner negativity) under dissipation. The dissipative regime introduces odd photon numbers into the SCLSs (see Figs. 3(c)-3(d)).
These states demonstrate significant squeezing in quadrature components: 3.37-3.37 dB ($\Delta^{2}p_{1}=0.23$) for mode $\hat{a}_{1}$ and 3.76-3.76 dB ($\Delta^{2}x_{2}=0.21$) for mode $\hat{a}_{2}$, along with super-Poissonian statistics confirmed by Fano factor analysis. The super-Poissonian character ($FF\gg 1$) enhances correlation-based imaging through photon bunching [a2017], making these states valuable for quantum sensing [Salykina2023, Singh2024] applications, particularly in Mach-Zehnder interferometry [Shukla2023, Toscano2006, Salykina2023, Singh2024] and optical qubit encoding [Yamamoto2020, Lloyd1999].
Odd SCLSs can be generated by changing the initial state to $|\psi_{0}\rangle=|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$, producing only odd photon numbers in mode $\hat{a}_{1}$: $e^{-itg\hat{H}_{int}}|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}\approx|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}-i2tg\alpha_{20}\sqrt{6}|3\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$. The analysis for odd SCLSs follows the same methodology as for even states and is omitted here. Recent attempts to find optimal generation timing for SCLSs remain an active research focus [Gorshenin2024].
A key finding is the formation of entanglement between the modes, a direct consequence of the fully quantized treatment of the pump. This entanglement, combined with the non-Gaussian characteristics of the SCLSs, creates a powerful resource for quantum technologies. Furthermore, the relationship between the observed non-classicality and Bell-type non-classicality, which generally differ in nature [Volovich2016], presents an interesting avenue for further investigation.
These states are particularly promising for enhancing the sensitivity of Mach-Zehnder interferometers [Shukla2023, Toscano2006, Salykina2023, Singh2024] and for applications in correlation-based imaging. From a practical perspective, these entangled non-Gaussian states can be generated in standard SPDC setups using periodically poled nonlinear crystals [see Belinsky2018, and references therein], requiring approximately 64% pump conversion efficiency. Their verification via homodyne tomography [Paul2024] is feasible and offers an advantage over $\chi^{(3)}$-based schemes due to lower pump intensities.
Our findings not only contribute to the toolbox of continuous-variable quantum information processing [Holevo2019] but also open new avenues for exploring the interplay between entanglement, non-Gaussianity, and non-classicality in fundamental quantum optical processes.
Acknowledgements.
We thank Prof. Anatoly V. Masalov for discussions on obtained results and providing valuable inputs.